Suppose the monopolist faces the market demand function given by Q=144/p2

The AVC of the firm is given as AVC = Q ½ and the firm has a fixed cost of $ 5

a) determine equilibrium P&Q

b) determine the maximum profit


1
Expert's answer
2021-05-19T19:16:48-0400

To identify the profit, there is a need to set the equation: MR = MC

Given:

"Q = \frac{144}{P^2}, \space P^2 = \frac{144}{Q}"


Or "P=(\frac{144}{Q})^{\frac{1}{2}} ...........(i)"


Therefore, "P=(\frac{12}{\sqrt{Q}})"


Now TR = Price and Quantity sold

TR = PQ

"=\frac{12}{\sqrt Q}\times Q"


"=12\sqrt Q..........................(ii)"



Now, MR=

"\frac{\Delta(TR)}{\Delta Q}" [MR derivative of "\frac{TR}{\Delta Q}" ]


"=\frac{\Delta 12\sqrt Q}{\Delta Q}"


"=\frac{1}{2}(12Q)\\=\frac{6}{\Delta Q}.........................(iii)"


Given the total fixed cost of $5


Average variable cost (AVC)"=Q^{\frac{1}{2}}=\sqrt Q"

TC = TFC + TVC

TVC = (AVC)(Q)

Therefore, total cost  "TC=5+(\sqrt Q)(Q)"

"=5+Q^{\frac{1}{2}+1}\\=5+Q^{\frac{3}{2}}.......................(iv)"

with TC, Now we can compute MC by taking the derivation within respect to Q

Now MR = MC,

Therefore to find the profit-maximizing level of output

"\frac{6}{\sqrt Q}=\frac{3}{2}\sqrt Q\\or\space Q=\frac{12}{3}"


Q = 4units

by putting Q = 4

"P=\frac{12}{\sqrt Q}"


"P=\frac{12}{\sqrt 4}"


"P=\frac{12}{2}"


P = $6

Then profit is:

Profit = PQ - TC

"=(6)(4) - [5 + (4)^{3}{2}]"

"=24-13\\
= \$11"